OpenStudy (anonymous):

In ∆ABC, AC = 15 centimeters, m B = 68°, and m C = 24°. What is BC to two decimal places? A. 6.58 centimeters B. 9.88 centimeters C. 13.57 centimeters D. 16.17 E. 19025 centimeters

OpenStudy (campbell_st):

start by finding x... using angle sum of a triangle 

OpenStudy (anonymous):

88?

OpenStudy (campbell_st):

then use the law of sines  \[\frac{15}{\sin(68)} = \frac{BC}{\sin(x)}\] hope it helps